OpenStudy (anonymous):

y - 2x = 3 3x - 2y = 5 solve by substitution. help? :/

OpenStudy (anonymous):

so basically, in substitution you want to solve for one variable in one of the equations

OpenStudy (anonymous):

x=1 and y=5

OpenStudy (anonymous):

so for example, y - 2x = 3 becomes y = 2x+3. then substitute y in the second equation as 2x+3 so 3x - 2y = 5 becomes 3x - 2(2x+3) = 5

OpenStudy (anonymous):

Solve for x. Then plug in the answer from x to find y. You should get anirud's answer

OpenStudy (anonymous):

Multiply The Equation 2x-y=3 with 2 and Assume it as Equation One and Take 3x-2y=5 as Equation 2 and Subtract 1-2 And Eliminate The y term and then U will Get The Value of x... Then Put the value of x in anyone of the Two Equations and U will get y

OpenStudy (anonymous):

You might want to make a note of this one: Whenever they ask you TO SOLVE WITH SUBSTITUTION: Either of the 2 equations, check which one looks simpler: y - 2x = 3 3x - 2y = 5 Wouldn't you say equation 1 looks more simple? :) So, just take the value y in terms of x: y = 3 + 2x. Then put the value of y in the second equation, so we have: 3x-2(3+2x) = 5 3x -6 -4x=5 -x -6=5 Take -x to the other side to get x, and bring 5 to the other side to get -5, -6-5=x -11=x. This is our first answer, let us substitute the value of x in equation 1. y-2(-11)=3 y+22=3 y=3-22 y=-19 So now we have the answer as: x=-11 and y=-19. Please let me know if you have any more doubts about this kind of question or anything else, I will try to explain as clearly as I can. :)